A 12.4 uF capacitor is connected through a 0.895-MΩ resistor to a constant potential difference of 60.0V.




a. Compute the charge at the following times after the connections are made: 5.0s,10.0s,20.0s,and 100.0s.




b. Compute the charging currents at the same instants.

1
Expert's answer
2022-03-22T15:47:54-0400

Answer

initial charge

Q'=CV="12.4*10^{-6}*60=0.744*10^{-3}C"

Rc=12.4*0.895=11.01sec


Chrge on capacitor

"Q=Q'(1-e^{\frac{t}{RC}})"

A) at the=5sec

Then charge

"Q=0.744*10^{-3}*(1-e^{\frac{5}{11.01}}) \\=0.28c"

Similirly we can find for other times.

B) current us given

"I=V\frac{e^{\frac{-t}{RC}}}{R*RC}"